OpenStudy (anonymous):

What is the slope of a line parallel to the line whose equation is y - x = 5? -1 1 5

OpenStudy (cwrw238):

rearranging y = x + 5 compare this with y = mx + c where m=slope of the line y = x + 5 y = mx + c m = 1 slope = 1

OpenStudy (anonymous):

First rearrange the equation into slope-intercept format (y=mx+b) so that you can identify the slope of the given line. Parallel lines always have the same slope.